Obtain expressions for reflection and transmission amplitude coefficients when
electric vector associated with a plane monochromatic e. m. wave is in the plane of
incidence.
1
Expert's answer
2019-08-23T11:09:45-0400

We note that any plane wave may be represented as a superposition of two orthogonal linearly polarized waves.

Ignoring the rapidly varying parts of the light wave and keeping only the complex amplitudes:

"E_i+E_r=E_t (1)"

where Ei is the incident amplitude, Er the reflected amplitude, Et is the transmitted amplitude


For the electric field the continuity relation becomes (θi= θr):

"-E_i×\cos θ_i + E_r×\cos θ_i = -E_t×\cos θ_t (2)"

We can write for the magnetic field

"-B_i×\cos θ_i + B_r×\cos θ_i = -B_t×\cos θ_t (3)"

The link between E and B is given by formula

"B=\frac {nE}{c} (4)"


We put (4) in (3)

"-n_i×(E_r- E_i)\cos θ_i = -n_t×E_t×\cos θ_t (5)"


We put (1) in (5)

"-n_i×(E_r- E_i)\cos θ_i = -n_t×(E_i+E_r)×\cos θ_t (5)"

Solving (5) for "\frac {E_r}{ E_i }" yields reflection and transmission coefficients for perpendicularly polarized light (Fresnel Equations)

"n=\frac {n_i×\cos θ_i -n_t×\cos θ_t}{n_i×\cos θ_i + n_t×\cos θ_t}"

"t=\frac {2n_i×\cos θ_i}{n_i×\cos θ_i + n_t×\cos θ_t}"